Jill has 24 coins in her pocket. How much money does she have?
1/2 are nickels
1/4 are quarters
1/6 are dimes
1/12 are pennies

Answer by stanbon(75887)

Jill has 24 coins in her pocket. How much money does she have?
1/2 are nickels = 12 nickels = 12*5 = 60 cents
1/4 are quarters = 6 quarters = 150 cents
1/6 are dimes = 4 dimes = 40 cents
1/12 are pennies = 2 pennies = 2 cents
----------------
Total = 60 + 150 + 40 + 2 = $2.52
========================
Cheers,
Stan H.
============